OpenStudy (anonymous):

A hotel lobby can seat 29 people at chairs and loveseats. The lobby has a total of 23 pieces of furniture for seating. Assuming a chair seats 1 person and a loveseat seats 2, how many loveseats are in the lobby? 4 5 6 or 7?

OpenStudy (anonymous):

the problem set up is C+L=29 1C+2L=23

OpenStudy (anonymous):

so 11?

OpenStudy (anonymous):

wait i think my set up was wrong

OpenStudy (anonymous):

yep it was 23 and 29 should be switched

OpenStudy (anonymous):

but wouldnt there have to be more than 7 loveseats?

OpenStudy (anonymous):

ummm no there are 6. C+L=23 -C-L=-23 C+2L=29 C+2L=29 let multply by -1 wen yu add this together yu get...?

OpenStudy (anonymous):

......that make no sence to me

OpenStudy (anonymous):

ummm ok chairs is C and loveseats is L get tht part?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

there are 23 pieces all together right?

OpenStudy (anonymous):

nvm i get it naw..thx

OpenStudy (anonymous):

welcome anytime